I am trying to find the solutions to:
(x^2-3)^2-43(x^2-3)+42=0
I can not figure out how to turn it into a quadratic and solve it.

Answer by lwsshak3(11628)

(x^2-3)^2-43(x^2-3)+42=0
let u=(x^2-3)
u^2-43u+42=0
(u-42)(u-1)=0
u=1=x^2-3
x^2=4
x=±√4=±2
or
x=42=x^2-3
x^2=45
x=±√45
solution: x=-√45, -2, 2, √45